A one billion(10^9) electrons flow over a period of 1µs from A to B in the wire . Determine current  flow through the wire in microampier.

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Current is the amount of charge flowing per unit time.

HenceI=\dfrac{q}{t}=\dfrac{ne}{t}=\dfrac{10^9\times 1.6\times 10^{-19}}{10^{-6}}A

=160\mu A